# EvaNet: Elevation-Guided Flood Extent Mapping on Earth Imagery

### v2

@@ A.1 Related Work on KGML
Two works utilize Fourier Neural Operator (FNO) to emulate (but speed up) hydrodynamic modeling/simulation for flood forecasting: [?] uses CREST-iMAP to generate training data on historical storm events where inputs to FNO include antecedent precipitation, simulated water depths, and DEM; while [?] builds an FNO-based surrogate for NEMO to produce accelerated emulation of coastal dynamics. Alternatives to FNO (to speed up PDE evaluation) for flood forecasting have been studied such as PINN [?]. There are also works that use LSTM to capture the temporal process, e.g., for streamflow prediction [?] and rainfall-runoff modeling [?].
@@ A.4 Model Setting Tuning
Effect of Fusion Operations. Recall from Figure 4 that we use 8 fusion operations (that generate the fused purple tensors). Table 5 now shows the performance of our model on test region R4, as well as its two variants where the 8 GLU operations are replaced with alternative fusion operations GIF [?] and GMU [?], respectively. We do not observe clear advantages of applying GIF and GMU over GLU in Figure 4, and the results are similar for the other test regions. Note that GLU is now widely adopted in recent works such as [???], rather than other fusion methods like GIF and GMU.
@@ A.6 Ablation Study with FCN-8s
Recall from Table 4 that when U-Net is used as the encoder-decoder backbone, both Eva-Reg and Eva-Conv improve the model performance. We remark that this improvement carries over to other ConvNet architectures. For example, Table 11 shows the ablation study results when the encoder-decoder backbone is FCN-8s from [?] instead of U-Net, where we observe similar conclusions.
@@ A.7 Effect as a U-Net Drop-in Replacement
Here we answer the following question: can EvaNet improve the performance of an existing model by replacing its U-Net component? Recall from Section 2 that the HMCT-PP (i.e., U-Net + HMCT) model [?] acts as a postprocessor of U-Net to further improve performance. If we replace U-Net with EvaNet, we hope that the HMCT postprocessor will provide an even better performance. Our experiments provides an affirmative answer. For example, Table 12 shows the performance of U-Net 7C and EvaNet 7C on test region R4, w/ and w/o HMCT postprocessing, respectively. We can see that ‘EvaNet 7C + HMCT’ gives the best performance, outperforming ‘U-Net 7C + HMCT.’

### v3

@@ A.1 Related Work on KGML
Two works utilize Fourier Neural Operator (FNO) to emulate (but speed up) hydrodynamic modeling/simulation for flood forecasting: Sun et al. [2023] uses CREST-iMAP to generate training data on historical storm events where inputs to FNO include antecedent precipitation, simulated water depths, and DEM; while Jiang et al. [2021] builds an FNO-based surrogate for NEMO to produce accelerated emulation of coastal dynamics. Alternatives to FNO (to speed up PDE evaluation) for flood forecasting have been studied such as PINN Mahesh et al. [2022]. There are also works that use LSTM to capture the temporal process, e.g., for streamflow prediction Ghosh et al. [2022] and rainfall-runoff modeling Nearing et al. [2020].
@@ A.4 Model Setting Tuning
Effect of Fusion Operations. Recall from Figure 4 that we use 8 fusion operations (that generate the fused purple tensors). Table 5 now shows the performance of our model on test region R4, as well as its two variants where the 8 GLU operations are replaced with alternative fusion operations GIF Arevalo et al. [2017] and GMU Kim et al. [2018], respectively. We do not observe clear advantages of applying GIF and GMU over GLU in Figure 4, and the results are similar for the other test regions. Note that GLU is now widely adopted in recent works such as Wu et al. [2019]; Hui et al. [2021b, a], rather than other fusion methods like GIF and GMU.
@@ A.6 Ablation Study with FCN-8s
Recall from Table 4 that when U-Net is used as the encoder-decoder backbone, both Eva-Reg and Eva-Conv improve the model performance. We remark that this improvement carries over to other ConvNet architectures. For example, Table 11 shows the ablation study results when the encoder-decoder backbone is FCN-8s from Long et al. [2015b] instead of U-Net, where we observe similar conclusions.
@@ A.7 Effect as a U-Net Drop-in Replacement
Here we answer the following question: can EvaNet improve the performance of an existing model by replacing its U-Net component? Recall from Section 2 that the HMCT-PP (i.e., U-Net + HMCT) model Sainju et al. [2022] acts as a postprocessor of U-Net to further improve performance. If we replace U-Net with EvaNet, we hope that the HMCT postprocessor will provide an even better performance. Our experiments provides an affirmative answer. For example, Table 12 shows the performance of U-Net 7C and EvaNet 7C on test region R4, w/ and w/o HMCT postprocessing, respectively. We can see that ‘EvaNet 7C + HMCT’ gives the best performance, outperforming ‘U-Net 7C + HMCT.’